# Forecasting realized volatility in the stock market: a path-dependent perspective

### v1

@@ Appendix B Out-of-sample forecast loss value and MCS test
GARCH $1.687\times 10^{-8}$ $8.075\times 10^{-5}$ $8.872$ $1.842$ $-8.379$
EGARCH $1.226\times 10^{-8}$ $7.551\times 10^{-5}$ $7.940$ $1.822$ $-8.378$
FGARCH $1.927\times 10^{-8}$ $1.113\times 10^{-4}$ $15.470$ $2.887$ $-8.299$
TGARCH $1.082\times 10^{-8}$ $4.808\times 10^{-5}$ $0.314$ $0.473$ $-6.994$
HAR-RV $1.284\times 10^{-8}$ $5.694\times 10^{-5}$ $3.708$ $0.950$ $-8.437$
HAR-CJ $5.082\times 10^{-9}$ $3.311\times 10^{-5}$ $0.525$ $0.519$ $-8.710$
HAR-RS $4.413\times 10^{-9}$ $3.282\times 10^{-5}$ $0.846$ $0.578$ $-8.736$
HAR-REX $5.119\times 10^{-9}$ $3.396\times 10^{-5}$ $0.594$ $0.554$ $-8.707$
HAR-REQ $5.099\times 10^{-9}$ $3.308\times 10^{-5}$ $0.529$ $0.524$ $-8.708$
HAR-PD-RV $6.986\times 10^{-9}$ $3.955\times 10^{-5}$ $0.817$ $0.578$ $-8.614$
HAR-PD-CJ $5.062\times 10^{-9}$ $3.204\times 10^{-5}$ $0.446$ $0.474$ $-8.710$
HAR-PD-RS $\mathbf{4.213\times 10^{-9}}$ $3.422\times 10^{-5}$ $1.059$ $0.630$ $-8.669$
HAR-PD-REX $4.964\times 10^{-9}$ $3.253\times 10^{-5}$ $0.525$ $0.510$ $-8.711$
LASSO-HAR-PD-CJ $5.068\times 10^{-9}$ $3.232\times 10^{-5}$ $0.464$ $0.483$ ${-8.711}$
LASSO-HAR-PD-RS $\underline{4.377\times 10^{-9}}$ $\underline{\mathbf{3.097\times 10^{-5}}}$ $0.677$ $0.491$ $-8.744$
LASSO-HAR-PD-REX $5.020\times 10^{-9}$ $3.277\times 10^{-5}$ $0.535$ $0.519$ $-8.709$
LASSO-HAR-PD-REQ $5.081\times 10^{-9}$ $3.438\times 10^{-5}$ $0.632$ $0.574$ $-8.702$
@@
Table 16: MCS test for each volatility model of CSI 300 with a forecast sample of 300-day
Model $H=1$ $H=5$ $H=22$
@@
HAR-CJ $1.000$ $0.991$ $1.000$ $0.985$ $0.000$ $0.000$ $0.990$ $0.559$ $0.000$ $0.000$ $0.993$ $0.559$
HAR-RS $1.000$ $0.991$ $1.000$ $1.000$ $1.000$ $0.002$ $0.972$ $0.359$ $0.000$ $0.000$ $0.972$ $0.359$
HAR-REX $0.000$ $0.000$ $1.000$ $0.448$ $0.000$ $0.000$ $0.816$ $0.481$ $0.000$ $0.000$ $0.816$ $0.481$
HAR-REQ $1.000$ $1.000$ $1.000$ $0.985$ $0.123$ $0.513$ $1.000$ $0.878$ $0.000$ $0.000$ $1.000$ $0.877$
HAR-PD-RV $0.000$ $0.000$ $0.974$ $0.929$ $0.000$ $0.000$ $0.000$ $0.000$ $0.000$ $1.000$ $0.851$ $0.000$
HAR-PD-CJ $0.392$ $0.540$ $1.000$ $1.000$ $0.303$ $0.548$ $1.000$ $0.672$ $0.000$ $0.000$ $1.000$ $0.652$
HAR-PD-RS $\mathbf{1.000}$ $\mathbf{1.000}$ $0.989$ $0.985$ $0.797$ $0.003$ $1.000$ $0.997$ $0.000$ $0.000$ $1.000$ $0.997$
HAR-PD-REX $1.000$ $1.000$ $1.000$ $0.985$ $0.995$ $\mathbf{1.000}$ $\mathbf{1.000}$ $\mathbf{1.000}$ $0.000$ $0.000$ $\mathbf{1.000}$ $\mathbf{1.000}$
@@
LASSO-HAR-PD-CJ $0.317$ $0.502$ $\mathbf{1.000}$ $0.191$ $0.259$ $0.513$ $0.996$ $0.615$ $0.000$ $0.000$ $0.995$ $0.978$
LASSO-HAR-PD-RS $\mathbf{1.000}$ $\mathbf{1.000}$ $0.997$ $0.864$ $\mathbf{1.000}$ $1.000$ $1.000$ $0.978$ $\mathbf{1.000}$ $\mathbf{1.000}$ $0.993$ $0.978$
LASSO-HAR-PD-REX $0.370$ $0.557$ $1.000$ $0.864$ $0.604$ $0.548$ $1.000$ $1.000$ $0.000$ $0.000$ $1.000$ $0.971$
LASSO-HAR-PD-REQ $1.000$ $0.927$ $\mathbf{1.000}$ $\mathbf{1.000}$ $0.000$ $0.000$ $0.612$ $0.621$ $0.000$ $0.000$ $0.612$ $0.621$
@@
Table 17: Out-of-sample forecast loss value of CSI 300 for the future 5-step
Model MSE MAE HMSE HMAE QLIKE
GARCH $1.201\times 10^{-8}$ $5.525\times 10^{-5}$ $2.226$ $0.905$ $-8.428$
EGARCH $1.806\times 10^{-8}$ $7.424\times 10^{-5}$ $4.124$ $1.348$ $-8.410$
FGARCH $2.555\times 10^{-8}$ $1.236\times 10^{-4}$ $10.718$ $2.559$ $-8.207$
TGARCH $1.065\times 10^{-8}$ $4.545\times 10^{-5}$ $0.873$ $0.680$ $-8.173$
HAR-RV $1.286\times 10^{-8}$ $5.516\times 10^{-5}$ $3.669$ $0.878$ $-8.367$
HAR-CJ $9.018\times 10^{-9}$ $3.846\times 10^{-5}$ $0.476$ $0.497$ $-8.543$
HAR-RS $7.581\times 10^{-9}$ $\mathbf{3.795\times 10^{-5}}$ $0.832$ $0.552$ $-8.514$
HAR-REX $9.645\times 10^{-9}$ $4.461\times 10^{-5}$ $0.807$ $0.671$ $-8.519$
HAR-REQ $9.352\times 10^{-9}$ $\underline{3.840\times 10^{-5}}$ $\underline{\mathbf{0.433}}$ $\underline{\mathbf{0.485}}$ $-8.531$
HAR-PD-RV $9.985\times 10^{-9}$ $4.602\times 10^{-5}$ $1.257$ $0.736$ $-8.507$
HAR-PD-CJ $8.873\times 10^{-9}$ $4.258\times 10^{-5}$ $0.696$ $0.611$ $-8.533$
HAR-PD-RS $\underline{\mathbf{7.575\times 10^{-9}}}$ $3.907\times 10^{-5}$ $0.948$ $0.583$ $-8.541$
HAR-PD-REX $8.502\times 10^{-9}$ $4.152\times 10^{-5}$ $0.662$ $0.607$ $-8.540$
LASSO-HAR-PD-CJ $8.878\times 10^{-9}$ $4.282\times 10^{-5}$ $0.712$ $0.619$ $-8.533$
LASSO-HAR-PD-RS $7.576\times 10^{-9}$ $3.907\times 10^{-5}$ $0.948$ $0.581$ $-8.542$
LASSO-HAR-PD-REX $8.745\times 10^{-9}$ $4.247\times 10^{-5}$ $0.682$ $0.618$ $-8.531$
LASSO-HAR-PD-REQ $8.182\times 10^{-9}$ $4.034\times 10^{-5}$ $0.629$ $0.583$ $-8.558$
  • 1.
Table 18: Out-of-sample forecast loss value of CSI 300 for the future 22-step
Model MSE MAE HMSE HMAE QLIKE
GARCH $1.254\times 10^{-8}$ $5.804\times 10^{-5}$ $1.002$ $1.022$ $-8.449$
EGARCH $2.051\times 10^{-8}$ $7.835\times 10^{-5}$ $8.450$ $1.641$ $-8.417$
FGARCH $7.790\times 10^{-8}$ $2.627\times 10^{-4}$ $5.054$ $1.369$ $-8.414$
TGARCH $1.071\times 10^{-8}$ $4.048\times 10^{-5}$ $0.656$ $0.589$ $-8.178$
HAR-RV $1.264\times 10^{-8}$ $5.235\times 10^{-5}$ $3.818$ $0.886$ $-8.417$
HAR-CJ $9.098\times 10^{-9}$ $3.647\times 10^{-5}$ $0.483$ $0.498$ $-8.590$
HAR-RS $7.613\times 10^{-9}$ $3.638\times 10^{-5}$ $0.814$ $0.553$ $-8.575$
HAR-REX $9.566\times 10^{-9}$ $4.221\times 10^{-5}$ $0.800$ $0.668$ $-8.567$
HAR-REQ $9.380\times 10^{-9}$ $\underline{\mathbf{3.599\times 10^{-5}}}$ $\underline{\mathbf{0.434}}$ $\underline{\mathbf{0.482}}$ $-8.583$
HAR-PD-RV $9.664\times 10^{-9}$ $4.243\times 10^{-5}$ $1.269$ $0.732$ $-8.557$
HAR-PD-CJ $9.071\times 10^{-9}$ $4.154\times 10^{-5}$ $0.766$ $0.635$ $-8.577$
HAR-PD-RS $\underline{\mathbf{7.625\times 10^{-9}}}$ $3.705\times 10^{-5}$ $0.941$ $0.574$ $-8.599$
HAR-PD-REX $8.751\times 10^{-9}$ $4.087\times 10^{-5}$ $0.716$ $0.629$ $-8.592$
LASSO-HAR-PD-CJ $9.075\times 10^{-9}$ $4.177\times 10^{-5}$ $0.782$ $0.643$ $-8.577$
LASSO-HAR-PD-RS $8.176\times 10^{-9}$ $3.687\times 10^{-5}$ $0.724$ $0.559$ $-8.597$
LASSO-HAR-PD-REX $9.124\times 10^{-9}$ $4.334\times 10^{-5}$ $0.790$ $0.667$ $-8.570$
LASSO-HAR-PD-REQ $8.556\times 10^{-9}$ $4.080\times 10^{-5}$ $0.750$ $0.625$ $-8.600$
  • 1.

### v2

@@ Appendix B Out-of-sample forecast loss value and MCS test
GARCH $1.128\times 10^{-8}$ $4.319\times 10^{-5}$ $0.539$ $0.501$ 8.301-8.301
EGARCH $1.578\times 10^{-8}$ $7.779\times 10^{-5}$ $4.720$ $1.440$ 8.115-8.115
TGARCH $1.071\times 10^{-8}$ $4.048\times 10^{-5}$ $0.656$ $0.589$ 8.178-8.178
GJRGARCH $1.138\times 10^{-8}$ $4.336\times 10^{-5}$ $0.506$ $0.493$ 8.269-8.269
HAR-RV $1.264\times 10^{-8}$ $5.235\times 10^{-5}$ $3.818$ $0.886$ 8.417-8.417
HAR-CJ $9.098\times 10^{-9}$ $3.647\times 10^{-5}$ $0.483$ $0.498$ 8.590-8.590
HAR-RS $7.613\times 10^{-9}$ $3.638\times 10^{-5}$ $0.814$ $0.553$ 8.575-8.575
HAR-REX $9.566\times 10^{-9}$ $4.221\times 10^{-5}$ $0.800$ $0.668$ 8.567-8.567
HAR-REQ $9.380\times 10^{-9}$ $\underline{\mathbf{3.599\times 10^{-5}}}$ $\underline{\mathbf{0.434}}$ $\underline{\mathbf{0.482}}$ 8.583-8.583
HAR-PD-RV $9.664\times 10^{-9}$ $4.243\times 10^{-5}$ $1.269$ $0.732$ 8.557-8.557
HAR-PD-CJ $9.071\times 10^{-9}$ $4.154\times 10^{-5}$ $0.766$ $0.635$ 8.577-8.577
HAR-PD-RS $\underline{\mathbf{7.625\times 10^{-9}}}$ $3.705\times 10^{-5}$ $0.941$ $0.574$ 8.599-8.599
HAR-PD-REX $8.751\times 10^{-9}$ $4.087\times 10^{-5}$ $0.716$ $0.629$ 8.592-8.592
LASSO-HAR-PD-CJ $9.075\times 10^{-9}$ $4.177\times 10^{-5}$ $0.782$ $0.643$ 8.577-8.577
LASSO-HAR-PD-RS $8.176\times 10^{-9}$ $3.687\times 10^{-5}$ $0.724$ $0.559$ 8.597-8.597
LASSO-HAR-PD-REX $9.124\times 10^{-9}$ $4.334\times 10^{-5}$ $0.790$ $0.667$ 8.570-8.570
LASSO-HAR-PD-REQ $8.556\times 10^{-9}$ $4.080\times 10^{-5}$ $0.750$ $0.625$ 8.600-8.600
@@
Table 18: MCS test for each volatility model of CSI 300 with a out-of-sample of 300-day
Model One-step-ahead Five-step-ahead Twenty-two-step-ahead
@@
HAR-CJ $1.000$ $0.986$ $1.000$ $0.982$ $1.000$ $1.000$ $0.990$ $0.559$ $0.992$ $0.998$ $0.993$ $0.999$
HAR-RS $1.000$ $0.998$ $0.987$ $1.000$ $1.000$ $1.000$ $0.972$ $0.359$ $1.000$ $0.998$ $0.972$ $1.000$
HAR-REX $0.000$ $0.000$ $0.934$ $0.458$ $0.791$ $0.427$ $0.816$ $0.481$ $0.984$ $0.432$ $0.816$ $0.481$
HAR-REQ $1.000$ $0.997$ $1.000$ $0.982$ $1.000$ $1.000$ $1.000$ $0.878$ $\mathbf{1.000}$ $\mathbf{1.000}$ $1.000$ $0.877$
HAR-PD-RV $0.000$ $0.000$ $0.819$ $0.911$ $0.000$ $0.673$ $0.640$ $0.000$ $0.993$ $0.998$ $0.851$ $0.000$
HAR-PD-CJ $0.365$ $0.546$ $1.000$ $0.755$ $0.303$ $0.894$ $0.848$ $0.672$ $0.963$ $0.993$ $0.611$ $0.417$
HAR-PD-RS $\mathbf{1.000}$ $\mathbf{1.000}$ $0.981$ $0.982$ $1.000$ $0.979$ $1.000$ $0.997$ $1.000$ $1.000$ $1.000$ $0.997$
HAR-PD-REX $0.796$ $0.816$ $1.000$ $0.985$ $0.995$ $\mathbf{1.000}$ $\mathbf{1.000}$ $\mathbf{1.000}$ $1.000$ $1.000$ 0.985{0.985} 0.871{0.871}
@@
LASSO-HAR-PD-CJ $0.296$ $0.505$ $\mathbf{1.000}$ $0.191$ $0.829$ $0.811$ $0.996$ $0.615$ $0.937$ $0.998$ $0.995$ $0.978$
LASSO-HAR-PD-RS $\mathbf{1.000}$ $\mathbf{1.000}$ $0.927$ $0.826$ $\mathbf{1.000}$ $1.000$ $1.000$ $0.978$ 1.000{1.000} 1.000{1.000} $\mathbf{1.000}$ $\mathbf{1.000}$
LASSO-HAR-PD-REX $0.366$ $0.504$ $1.000$ $0.755$ $0.715$ $0.640$ $1.000$ $1.000$ $0.499$ $0.411$ $1.000$ $0.971$
LASSO-HAR-PD-REQ $1.000$ $0.927$ $\mathbf{1.000}$ $\mathbf{1.000}$ $1.000$ $0.979$ $0.612$ $0.621$ $1.000$ $0.998$ $0.612$ $0.621$